OpenStudy (anonymous):

Find the equation in form f(x)=a(x-h)^2+k of the parabola that has vertex (-8,-11) at and passes through the point (-10,9)

OpenStudy (anonymous):

f(x)=a(x+8)^2-11 passes the point (-10,9) 9=a[(-10)+8]^2-11 9+11=-2a a=-10 Rewrite f(x)= -10(x+8)^2-11

OpenStudy (campbell_st):

umm isn't there a slight error in the value of a 9 + 11 = a(-2)^2 a = 5 so the equation is y = 5(x + 8)^2 - 11

OpenStudy (campbell_st):

oops forgot the line 20 = 4a so a = 5

OpenStudy (campbell_st):

and that make sense as (-8, -11) is the lowest point and (-10, 9) is above it

OpenStudy (anonymous):

Oh yes, thank you @campbell_st!!